Suppose a sample of 30 JJC students are given an IQ test. If the sample has...

  1. Suppose a sample of 30 JJC students are given an IQ test. If the sample has a standard deviation of 12.23 points, find a 90% confidence interval for the population standard deviation.
Answer #1

Solution :

Given that,

s = 12.23

s2 = 149.5729

\chi2R = \chi 2\alpha/2,df = 42.557

\chi2L = \chi 21 - \alpha /2,df = 17.708

The 90% confidence interval for \sigma is,

\sqrt(n - 1)s2 / \chi 2\alpha/2 < \sigma < \sqrt (n - 1)s2 / \chi 21 - \alpha /2

\sqrt(29)(149.5729) / 42.557 < \sigma < \sqrt (29)(149.5729) / 17.708

10.10 < \sigma < 15.65

(10.10 , 15.65)